OpenStudy (anonymous):

Find the ratio of x to y. 3x = 8y

OpenStudy (anonymous):

x = 8/3y

OpenStudy (mathmale):

I believe that "find the ratio of x to y" literally means solving for (x/y). Starting with FutureMathProfessor's result, x = 8/3y, simply divide both sides of the equation by y: \[x=\left( 8/3 \right)y\] becomes \[\frac{ x }{ y }=\frac{ 8 }{3}.\] That's it! In other words, that's the "ratio of x to y."

OpenStudy (anonymous):

3x means it is 3 times x. 8y means it is 8 times y. 3 times x equals 8 times y. That means, there needs to be 8/3 times y for each x.